Why can osmometry be used to determine the molar mass of a solute when colligative properties are, by definition, independent of the identity of the solute.
Answer #1

Colligative properties are those properties of dilute solutions which depend only on the number of particles of the solute but independent of it's nature.

Osmotic pressure or osmometry is mainly used to find the molar mass of the solute.

For dilute solutions, osmotic pressure is directly proportional to the molarity (C) of the solution at a given temperature (T).

\pi = C R T = (n/V) R T

where, \pi – osmotic pressure, C – molarity, R – gas constant, T – temperature, V – a volume of solution in liters, n – the number of moles of solute.

Also,n=(weight of solute in the solvent)/(molar mass of solute).= w/M

Thus,\pi= (w/MV) RT.

Rearranging the above equation, we get, M = (w/\piV) RT. This gives us the molar mass of the solute.